OpenStudy (anonymous):

Sin(3x-pi/2) ALGEBRAICALLY FIND POINTS

OpenStudy (anonymous):

so where did we end at.... i think it was \[\frac{\pi}{2}+2\pi n\]

OpenStudy (anonymous):

for when Sin(u) = 1

OpenStudy (lilsis76):

yes and u told me N is any number

OpenStudy (anonymous):

yep, alright so if we know that \[u=\frac{\pi}{2}+2\pi n\] to make sin(u)=0 \[sin(\frac{\pi}{2}+2\pi n)=1\]

OpenStudy (anonymous):

just to make sure lets test it out... for n=1 we get \[sin(\frac{\pi}{2}+2\pi)=sin(\frac{5\pi}{2})=\] since \[\frac{4\pi}{2}=2\pi\] starting from zero, we go around the circle once and then go another \[\frac{\pi}{2}\] which is the same angle as \[\frac{\pi}{2}\] sin at that point = 1! so we know that it's correct!

OpenStudy (anonymous):

now remember how we made \[u=3x-\frac{pi}{2}\]

OpenStudy (anonymous):

?

OpenStudy (lilsis76):

hmmm... hold on let me see again

OpenStudy (anonymous):

or we can look at it this way we know that \[sin(\frac{\pi}{2}+2\pi n)=1\] if we want \[sin(3x-\frac{\pi}{2})\] to equal 1 also we should equal them to each other if we do that we get \[sin(\frac{\pi}{2}+2\pi n)=sin(3x-\frac{\pi}{2})\]

OpenStudy (lilsis76):

im trying to see how u got 5pi/2 hold on im trying to write it out

OpenStudy (anonymous):

alright =]

OpenStudy (anonymous):

let me kwow when you're done with that =]

OpenStudy (lilsis76):

okay i got it, i see how u got the 5pi/2

OpenStudy (anonymous):

alright does what i said after make sense?

OpenStudy (lilsis76):

ya, the main problem was y=3x - pi/2 and you said that sin(pi/2 +2pi n) and that one is from---^ is that to get a point? i forgot already. WHY DOES THE INTERNET KEEP DOING THAT?!

OpenStudy (anonymous):

sin(pi/2+2pi n) is for when sin(u)=1

OpenStudy (lilsis76):

okay i got that now then

OpenStudy (anonymous):

so sin(pi/2+2pi n) will always be 1 no matter what integer you let n be

OpenStudy (lilsis76):

okay, i wrote that down and circled it

OpenStudy (anonymous):

no you want that to be the same for your problem \[sin(3x-\frac{\pi}{2})\] in other words if you let \[sin(3x-\frac{\pi}{2})=1\] and \[sin(\frac{\pi}{2}+2\pi n)=1 \] they'd be equal to each other correct?

OpenStudy (anonymous):

i mean't now* not no lol =]

OpenStudy (anonymous):

cause \[1=1\]

OpenStudy (lilsis76):

haha ya i understand that

OpenStudy (anonymous):

so now you can make them equal to each other \[sin(3x-\frac{\pi}{2})=sin(\frac{\pi}{2}+\pi n)\]

OpenStudy (anonymous):

the sins will cancel correct?

OpenStudy (lilsis76):